OpenStudy (ksaimouli):

How do you rearrange the equation (2x)/(4pi)+((1-x)/2)=0 to solve for x?

mathslover (mathslover):

\[\huge{(2x)\pi+\frac{1-x}{2}=0}\] \[\huge{\frac{4x\pi +1-x}{2}=0}\]

OpenStudy (anonymous):

That's what I was looking for but I don't know how to type that in.

mathslover (mathslover):

\[\huge{4x\pi+1-x=0}\] \[\huge{4x\pi-x=-1}\] \[\huge{x(4\pi-1)=-1}\]

OpenStudy (ksaimouli):

@mathslover how did u get 1 step

mathslover (mathslover):

oops i didnot see that question clearly

OpenStudy (ksaimouli):